OpenStudy (anonymous):

find all the real values of x which satisfy x^2-3x+2>0 and x^2-3x-4<= 0 (now both the inequalities will give different solutions, will the solution of one has to satisfy the other inequality, pls explain)

OpenStudy (anonymous):

did you solve \(x^2-3x+2>0\) ?

OpenStudy (anonymous):

the key word in the question "find all the real values of x which satisfy \(x^2-3x+2>0\) AND \( x^2-3x-4\leq 0\)" is AND solve each separately and take the intersection of the two sets

OpenStudy (anonymous):

now the question is, do you know the solution to \(x^2-3x+2>0\) ?

OpenStudy (anonymous):

yes i did manage to factorize both x^2−3x+2=0 AND x^2−3x−4=0 .. the insertion part I am not getting.

OpenStudy (anonymous):

sorry it will be the intersection and not insertion

OpenStudy (anonymous):

it is more than just factoring once you have \(x^2-3x+2=(x-1)(x-2)\) you see that it is positive on the intervals \[(-\infty,1)\cup (2,\infty)\]

OpenStudy (anonymous):

is that part clear?

OpenStudy (anonymous):

yes how about the other part and combining both ?

OpenStudy (anonymous):

which means 1<x<2 in this interval x^2−3x+2 is positive

OpenStudy (anonymous):

and the solution to the next one is \((-1,4)\) if i am not mistaken

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

\(y=x^2-3x+2\) is a parabola that opens UP

OpenStudy (anonymous):

it is negative between the zeros, positive outside them

OpenStudy (anonymous):



OpenStudy (anonymous):

in the interval \((1,2)\) it is NEGATIVE you want positive

OpenStudy (anonymous):

ok but how to get a solution that satisfies both ... how to write the answer ?

OpenStudy (anonymous):

well we can't hope to get them together until we know what they are separately first !

OpenStudy (anonymous):

the solution to the first one is \(x<1\) or \(x>2\) better written in interval notation as \[(-\infty,1)\cup (2,\infty)\]

OpenStudy (anonymous):

yes i get u

OpenStudy (anonymous):

the solution to the second one is \(-1\leq x\leq 4\) which is the same as \[[-1,4]\]

OpenStudy (anonymous):

now since you are given an AND statement you need to find the intersection of the two sets; the values of \(x\) that satisfy both

OpenStudy (anonymous):



OpenStudy (anonymous):

what is in common to both of these sets? the interval \([-1,1)\) is in both of them for example

OpenStudy (anonymous):

so it will be between -1 and 1 and 2 and 4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

be careful with the inequalities because one is \(\leq \) and the other is \(<\)

OpenStudy (anonymous):

-1<=x<1 and 2<x<= 4 ?

OpenStudy (anonymous):

in interval notation the one on the left is \([-1,1)\) or you can write \(-1\leq x<1\)

OpenStudy (anonymous):

yes that looks good

OpenStudy (anonymous):

although if you are a stickler for details, you should not use the word "and" in your answer

OpenStudy (anonymous):

hey many thanks ....u explained it extremely well :) ..

OpenStudy (anonymous):

-1<=x<1 OR 2<x<= 4

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

how do i give a medal ? i am new to open study ?